# NetMamba: Efficient Network Traffic Classification via Pre-training Unidirectional Mamba

### v3

@@ Abstract
Network traffic classification is a crucial research area aiming to enhance service quality, streamline network management, and bolster cybersecurity. To address the growing complexity of transmission encryption techniques, various machine learning and deep learning methods have been proposed. However, existing approaches face two main challenges. Firstly, they struggle with model inefficiency due to the quadratic complexity of the widely used Transformer architecture. Secondly, they suffer from inadequate traffic representation because of discarding important byte information while retaining unwanted biases. To address these challenges, we propose NetMamba, an efficient linear-time state space model equipped with a comprehensive traffic representation scheme. We adopt a specially selected and improved unidirectional Mamba architecture for the networking field, instead of the Transformer, to address efficiency issues. In addition, we design a traffic representation scheme to extract valid information from massive traffic data while removing biased information. Evaluation experiments on six public datasets encompassing three main classification tasks showcase NetMamba’s superior classification performance compared to state-of-the-art baselines. It achieves an accuracy rate of nearly 99% (some over 99%) in all tasks. Additionally, NetMamba demonstrates excellent efficiency, improving inference speed by up to 60 times while maintaining comparably low memory usage. Furthermore, NetMamba exhibits superior few-shot learning abilities, achieving better classification performance with fewer labeled data. To the best of our knowledge, NetMamba is the first model to tailor the Mamba architecture for networking.
@@ I Introduction
Specifically, NetMamba initially extracts hierarchical flow information from raw traffic and converts it into a stride sequence, which serves as the model’s input. Subsequently, NetMamba undergoes self-supervised pre-training on large unlabeled datasets using a masked autoencoder structure, which is designed to learn generic representations of traffic data through reconstructing masked strides. Finally, the decoder is replaced with a multi-layer perceptron head, and NetMamba is fine-tuned on limited labeled data to refine traffic representations and adapt to downstream traffic classification tasks. Extensive experiments conducted on publicly available datasets demonstrate the effectiveness and efficiency of NetMamba. In all classification tasks, NetMamba achieves an accuracy rate of nearly 99% (some over 99%), and, compared to existing baselines, it improves inference speed by up to 60 times while maintaining low GPU memory usage. Furthermore, NetMamba exhibits superior few-shot learning capabilities in comparison to other pre-training models, achieving better performance with fewer labeled data.
@@ II-C Traffic Representation Schemes
Recent advancements in deep learning have endeavored to utilize raw traffic bytes. However, as shown in Table I, these methods face limitations. They often neglect crucial information in packet headers and introduce unwanted biases by retaining biased IP addresses, ignoring byte balance, or using improper data-splitting techniques.
@@
Method Header Payload IA1 BB2 Splitting
@@
    IA: IP Anonymizing removes all IP addresses.
  • 2
    BB: Byte Balance sets fixed sizes for headers and payloads.
@@ III-B Discretization
where $\mathbf{\overline{A}}=\exp(\Delta\mathbf{A})$ and $\mathbf{\overline{B}}\approx\Delta\mathbf{B}$ represent the discretized parameters, with $\Delta$ denoting the discretization step size. This recurrent formulation is characterized by linear time complexity.
@@ V-2 Packet Parsing
For each flow, all packets are processed through several sequential operations to preserve valuable information and eliminate unnecessary interference. When narrowing down the scope for analyzing traffic data related to specific applications or services, we exclude all packets carried by non-IP protocols, such as Address Resolution Protocol (ARP) and Dynamic Host Configuration Protocol (DHCP). Considering the critical information contained within both the packet (e.g., the total length field) and the payload (text content for upper-level protocols), we choose to retain these elements. Furthermore, to mitigate biases introduced by identifiable information, all packets are anonymized through the removal of Ethernet headers and the masking of IP addresses.
@@ VII-A2 Comparison Methods
    Classical machine learning methods such as AppScaner [3] and FlowPrint [2] that rely on statistical features for traffic classification.
@@ VII-A4 Evaluation Metrics
We assess and compare the performance of NetMamba using four typical metrics: Accuracy(AC), Precision(PR), Recall(RC), and weighted F1 Score(F1).
@@ VII-B Overall Evaluation
ET-BERT[9] 187.4 136.4 0.9386 0.9451 0.9386 0.9401 0.9105 0.8809 0.9105 0.8850 0.9937 0.9938 0.9937 0.9937
YaTC(OF)[10] 2.3 2.1 0.9177 0.9203 0.9177 0.9176 0.9277 0.9277 0.9277 0.9262 0.9949 0.9949 0.9949 0.9949
YaTC [10] 2.3 2.1 0.9042 0.9081 0.9042 0.9042 0.9310 0.9307 0.9310 0.9295 0.9959 0.9959 0.0059 0.9959
NetMamba 2.2 1.9 0.9869 0.9871 0.9869 0.9864 0.9881 0.9885 0.9881 0.9881 0.9985 0.9985 0.9985 0.9985
@@
ET-BERT[9] 187.4 136.4 0.9980 0.9981 0.9980 0.9980 0.9566 0.9566 0.9566 0.9565 0.9910 0.9911 0.9910 0.9910
YaTC(OF)[10] 2.3 2.1 0.9986 0.9986 0.9986 0.9986 0.9848 0.9848 0.9848 0.9848 0.9951 0.9950 0.9951 0.9950
YaTC [10] 2.3 2.1 0.9959 0.9959 0.9959 0.9959 0.9819 0.9820 0.9819 0.9819 0.9947 0.9749 0.9747 0.9734
NetMamba 2.2 1.9 0.9993 0.9993 0.9993 0.9993 0.9899 0.9899 0.9899 0.9899 0.9990 0.9991 0.9990 0.9990
@@
We evaluated the performance of NetMamba in categorizing traffic using six publicly available datasets. As shown in Table IV and Table V, NetMamba demonstrates superior performance compared to all baseline methods across five datasets. However, it falls slightly short in comparison to TFE-GNN on the CICIoT2022 dataset, with a marginal 0.15 percentage point difference in accuracy and f1 score. On average, NetMamba achieves accuracy levels ranging from 0.9869 to 0.9993 and f1 scores between 0.9864 and 0.9993. Notably, NetMamba maintains the fewest parameters among all deep learning methods, underscoring its efficient yet effective capabilities in traffic representation learning.
VII-B1 CrossPlatform (Android)
As indicated in Table IV, NetMamba demonstrates significant improvements over existing methods on the CrossPlatform(Android) dataset. Specifically, our model achieves a notable improvement of 4.83% in accuracy and 4.64% in f1 score compared to the state-of-the-art method (ET-BERT). ET-BERT focuses solely on learning traffic representations from packet payloads, overlooking the valuable content information carried by packet headers. In contrast, NetMamba effectively models both header and payload characteristics, leading to a more comprehensive analysis of traffic patterns.
VII-B2 CrossPlatform (iOS)
On the CrossPlatform(iOS) dataset, NetMamba surpasses all baseline methods and outperforms the state-of-the-art technique (YaTC) by more than 5% across all evaluation metrics. Beyond the variances in base model architecture, NetMamba establishes a more resilient traffic representation scheme compared to YaTC. This enhancement is achieved through techniques such as IP masking and stride cutting, contributing to a stronger overall performance.
VII-B3 CICIoT2022
Concerning the CICIoT2022 dataset, NetMamba outperforms all baselines except TFE-GNN, which achieves a slight advantage of 0.15% across all evaluation metrics. TFE-GNN represents traffic flow through a byte-level correlation graph and utilize graph neural networks to capture traffic patterns. However, TFE-GNN lags significantly behind NetMamba on other datasets, particularly trailing by over 23% on the ISCXTor2016 dataset, indicating its unstable classification performance.
VII-B4 ISCXTor2016, ISCXVPN2016 & USTC-TFC2016
As depicted in Table VI, NetMamba surpasses all existing methods across the three encrypted datasets. Particularly noteworthy is the significantly unstable performance exhibited by methods other than YaTC and YaTC(OF) across different datasets. Given the similarities between NetMamba and YaTC, we contend that a robust traffic representation scheme, which incorporates both header and payload data, alongside a well-designed pre-training task, plays a crucial role in enhancing encrypted traffic analysis capabilities.
@@ VII-C Inference Efficiency Evaluation
In Figure 4(b), NetMamba demonstrates lower GPU memory consumption than most models, except FS-Net, YaTC and NT-Linear, when using large batch sizes. FS-Net’s reliance on RNNs, which require linear memory relative to sequence length, reduces memory costs but results in slower inference and poorer classification performance. YaTC reduces memory usage by shortening input sequence length through a model forward trick. Without such an optimization, YaTC(OF) consumes up to four times more GPU memory than NetMamba As shown in the subsequent ablation study, NT-Linear exhibits unstable classification performance due to over-compression of standard attention mechanisms. Compared to other baselines, NetMamba achieves improved memory efficiency primarily by customizing GPU operators that minimize the storage of extensive intermediate states and conduct recomputation during the backward pass.
@@
NetMamba (default) 0.9869 0.9869 0.9881 0.9881 0.9985 0.9985 0.9993 0.9993 0.9899 0.9899 0.9990 0.9990
w/ Bidirectional Mamba1 0.9852 0.9851 0.9793 0.9793 0.9974 0.9974 0.9986 0.9986 0.9935 0.9935 0.9987 0.9987
w/ Cascading Mamba1 0.8939 0.8942 0.9470 0.9455 0.9826 0.9826 0.9952 0.9952 0.9646 0.9643 0.9938 0.9937
w/ Vanilla Transformer2 0.9815 0.9815 0.9755 0.9755 0.9979 0.9980 0.9986 0.9986 0.9928 0.9927 0.9997 0.9997
w/ Linear Transformer2 0.9806 0.9820 0.9871 0.9871 0.8954 0.8998 0.9986 0.9986 0.8295 0.8398 0.4014 0.4226
w/o Position Embedding 0.9845 0.9841 0.9746 0.9726 0.9959 0.9959 1.0000 1.0000 0.9877 0.9877 0.9967 0.9966
w/o Pre-training 0.9807 0.9810 0.9779 0.9761 0.9949 0.9949 0.9959 0.9959 0.9798 0.9797 0.9984 0.9984
w/o Header 0.6933 0.7888 0.8430 0.8918 0.5525 0.5333 0.8464 0.8530 0.5441 0.4821 0.5136 0.5675
w/o Payload 0.9040 0.9040 0.9382 0.9383 0.9913 0.9913 1.0000 1.0000 0.9675 0.9676 0.9964 0.9964
w/o IP Masking 0.9242 0.9244 0.9403 0.9404 0.9959 0.9959 0.9979 0.9979 0.9812 0.9812 0.9964 0.9962
w/ Patch Splitting3 0.9826 0.9826 0.9784 0.9770 0.9979 0.9979 0.9993 0.9993 0.9892 0.9892 0.9974 0.9974
@@ VII-D1 Model-level Ablation
Initially, we replaced all unidirectional Mamba blocks in NetMamba with bidirectional ones used by [12]. The experimental results revealed a slight performance decline across all datasets except for ISCXVPN2016. This suggests that unidirectional Mamba is well-suited for processing network traffic data, given that packets are transmitted sequentially and earlier packets possess limited information about subsequent ones. Moreover, incorporating bidirectional or even omnidirectional Mamba blocks introduces additional computational and memory overheads due to extra scan passes, ultimately reducing efficiency. Thus, unidirectional Mamba stands out as the preferable choice.
@@
To clarify the source of NetMamba’s classification performance gains over existing Transformer-based models, we further evaluated two Transformer-based ablation variants, NT-Vanilla and NT-Linear. The results show that NT-Vanilla performs similarly to NetMamba across all datasets, suggesting that our proposed traffic representation scheme plays a key role in enhancing classification performance. With this representation scheme, a linear-time Mamba-based classifier achieves classification performance comparable to that of a quadratic-time Transformer-based model. However, NT-Linear, due to information loss from the overcompression of attention mechanisms, exhibits unstable performance and falls significantly behind NetMamba on three datasets.
While positional information is inherently preserved in sequence models such as Mamba, eliminating explicit positional embedding still results in a reduction in accuracy ranging from 0.03% to 1.35% across all datasets except for ISCXTor2016. This suggests that reinforced positional information aids the model in capturing correlations within sequential traffic data.
The pre-training process is designed to capture general traffic understanding from extensive unlabeled data. When compared to the non-pre-trained counterpart, pre-trained NetMamba demonstrates accuracy improvements ranging from 0.06% to 1.02%, affirming the effectiveness of our MAE-based pre-training task.

### v4

@@ Abstract
Network traffic classification is a crucial research area aiming to enhance service quality, streamline network management, and bolster cybersecurity. To address the growing complexity of transmission encryption techniques, various machine learning and deep learning methods have been proposed. However, existing approaches face two main challenges. Firstly, they struggle with model inefficiency due to the quadratic complexity of the widely used Transformer architecture. Secondly, they suffer from inadequate traffic representation because of discarding important byte information while retaining unwanted biases. To address these challenges, we propose NetMamba, an efficient linear-time state space model equipped with a comprehensive traffic representation scheme. We adopt a specially selected and improved unidirectional Mamba architecture for the networking field, instead of the Transformer, to address efficiency issues. In addition, we design a traffic representation scheme to extract valid information from massive traffic data while removing biased information. Evaluation experiments on six public datasets encompassing three main classification tasks showcase NetMamba’s superior classification performance compared to state-of-the-art baselines. It achieves accuracy rates exceeding 90%, with some surpassing 99%, across all tasks. Additionally, NetMamba demonstrates excellent efficiency, improving inference speed by up to 60 times while maintaining comparably low memory usage. Furthermore, NetMamba exhibits superior few-shot learning abilities, achieving better classification performance with fewer labeled data. To the best of our knowledge, NetMamba is the first model to tailor the Mamba architecture for networking.
@@ I Introduction
Specifically, NetMamba initially extracts hierarchical flow information from raw traffic and converts it into a stride sequence, which serves as the model’s input. Subsequently, NetMamba undergoes self-supervised pre-training on large unlabeled datasets using a masked autoencoder structure, which is designed to learn generic representations of traffic data through reconstructing masked strides. Finally, the decoder is replaced with a multi-layer perceptron head, and NetMamba is fine-tuned on limited labeled data to refine traffic representations and adapt to downstream traffic classification tasks. Extensive experiments conducted on publicly available datasets demonstrate the effectiveness and efficiency of NetMamba. In all classification tasks, NetMamba consistently achieves accuracy rates above 90%, with some exceeding 99%. Compared to existing baselines, it improves inference speed by up to 60 times while maintaining low GPU memory usage. Furthermore, NetMamba exhibits superior few-shot learning capabilities in comparison to other pre-training models, achieving better performance with fewer labeled data.
@@ II-C Traffic Representation Schemes
Recent advancements in deep learning have endeavored to utilize raw traffic bytes. However, as shown in Table I, these methods face limitations. They often neglect crucial information in packet headers and introduce unwanted biases by ignoring byte balance or using improper data-splitting techniques.
@@
Method Header Payload Byte Balance2 Splitting
@@
    Byte Balance sets fixed sizes for headers and payloads.
@@ III-B Discretization
where $\mathbf{\overline{A}}=\exp(\Delta\mathbf{A})$ and $\mathbf{\overline{B}}\approx\Delta\mathbf{B}$ represent the discretized parameters, with $\Delta$ denoting the discretization step size. This recurrent formulation, known for its linear time complexity, is suitable for model inference but lacks parallelizability during training.
By expanding Equation 2, SSMs can be transformed into convolutional formulations as follows:
$\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{L-1}\mathbf{\overline{B}})$ (3)
$y$ $=x*\mathbf{\overline{K}}$
where $L$ represents the length of the input sequence $x$, * denotes the convolution operation, and $\mathbf{\overline{K}}\in\mathbb{R}^{L}$ refers to a structured convolutional kernel. This convolutional representation solves the computational parallelization dilemma encountered in the recurrent version.
@@ V-2 Packet Parsing
For each flow, all packets are processed through several sequential operations to preserve valuable information and eliminate unnecessary interference. When narrowing down the scope for analyzing traffic data related to specific applications or services, we exclude all packets carried by non-IP protocols, such as Address Resolution Protocol (ARP) and Dynamic Host Configuration Protocol (DHCP). Considering the critical information contained within both the packet (e.g., the total length field) and the payload (text content for upper-level protocols), we choose to retain these elements. Furthermore, to mitigate biases introduced by identifiable information, all packets are anonymized through the removal of Ethernet headers.
@@ VII-A1 Datasets
We observed an imbalance in flow counts across traffic categories, which adversely impacts model performance. To address this, we set upper and lower flow limits for each category. Categories below the lower limit are discarded, while those above the upper limit are randomly sampled.
@@ VII-A2 Comparison Methods
    Classical machine learning methods such as AppScanner [3] and FlowPrint [2] that rely on statistical features for traffic classification.
@@ VII-A4 Evaluation Metrics
We assess the performance of NetMamba using four typical metrics: Accuracy(AC), Precision(PR), Recall(RC), and weighted F1 Score(F1).
@@ VII-B Overall Evaluation
ET-BERT[9] 187.4 136.4 0.8743 0.8913 0.8743 0.8786 0.9105 0.8809 0.9105 0.8850 0.9937 0.9938 0.9937 0.9937
YaTC(OF)[10] 2.3 2.1 0.9076 0.9107 0.9076 0.9077 0.9263 0.9282 0.9263 0.9264 0.9949 0.9949 0.9949 0.9949
YaTC [10] 2.3 2.1 0.8952 0.8989 0.8952 0.8952 0.9270 0.9296 0.9270 0.9272 0.9974 0.9975 0.9974 0.9974
NetMamba 2.2 1.9 0.9094 0.9133 0.9094 0.9096 0.9301 0.9327 0.9301 0.9305 0.9928 0.9931 0.9928 0.9929
@@
ET-BERT[9] 187.4 136.4 0.9967 0.9967 0.9967 0.9967 0.9566 0.9566 0.9566 0.9565 0.9910 0.9911 0.9910 0.9910
YaTC(OF)[10] 2.3 2.1 0.9986 0.9986 0.9986 0.9986 0.9805 0.9808 0.9805 0.9806 0.9960 0.9955 0.9960 0.9957
YaTC [10] 2.3 2.1 0.9959 0.9959 0.9959 0.9959 0.9848 0.9849 0.9848 0.9848 0.9972 0.9976 0.9972 0.9970
NetMamba 2.2 1.9 0.9986 0.9986 0.9986 0.9986 0.9805 0.9808 0.9805 0.9806 0.9960 0.9957 0.9960 0.9957
@@
We evaluated the performance of NetMamba in categorizing traffic using six publicly available datasets. As shown in Table IV and Table V, NetMamba consistently outperforms all baseline methods on three datasets and ranks second on two others. However, it falls slightly short on the CICIoT2022 dataset, with a maximum difference of 0.72 percentage points in both accuracy and F1 score. On average, NetMamba achieves accuracy levels between 0.9094 and 0.9986, and F1 scores ranging from 0.9096 to 0.9986. Notably, NetMamba maintains the fewest parameters among all deep learning methods, underscoring its efficient yet effective capabilities in traffic representation learning.
VII-B1 CrossPlatform (Android & iOS) and ISCXTor2016
The CrossPlatform Android and iOS datasets consist of encrypted traffic generated by the top 100 applications from the US, China, and India, encompassing over 200 categories in both cases. Additionally, ISCXTor2016 includes application traffic using the Onion Router (Tor) for encrypted communications.
As shown in Table IV, NetMamba demonstrates superior performance across these three datasets, with F1 score improvements ranging from 0.18% to 0.31%. In comparison to state-of-the-art pre-trained methods, ET-BERT exclusively focuses on learning traffic representations from packet payloads, overlooking the valuable information carried by packet headers. Moreover, beyond the variances in base model architecture, YaTC differs from NetMamba in the traffic representation scheme, employing a two-dimensional splitting technique. In contrast, NetMamba effectively models both header and payload characteristics and optimizes traffic data segmentation to mitigate biases. This leads to a more comprehensive analysis of traffic patterns.
VII-B2 CICIoT2022, ISCXVPN2016 & USTC-TFC2016
The CICIoT2022 dataset consists of traffic collected from a laboratory network designed for profiling, behavioral analysis, and vulnerability testing of various IoT devices. The ISCXVPN2016 dataset includes encrypted communication traffic tunneled through Virtual Private Networks (VPN). The USTC-TFC2016 dataset comprises encrypted traffic from both malware and benign applications.
As depicted in Table VI, NetMamba performs only slightly worse than YaTC on the ISCXVPN2016 and USTC-TFC2016 datasets, while achieving comparable performance to other state-of-the-art methods on the CICIoT2022 dataset. Although TFE-GNN achieves the highest performance on the CICIoT2022 dataset, this non-pre-trained model falls significantly behind NetMamba on the other datasets, highlighting its unstable classification performance.
@@ VII-C Inference Efficiency Evaluation
In Figure 4(b), NetMamba demonstrates lower GPU memory consumption than most models, except FS-Net, YaTC and NT-Linear, when using large batch sizes. FS-Net’s reliance on RNNs, which require linear memory relative to sequence length, reduces memory costs but results in slower inference and poorer classification performance. YaTC reduces memory usage by shortening input sequence length through a model forward trick. Without such an optimization, YaTC(OF) consumes up to four times more GPU memory than NetMamba. As shown in the subsequent ablation study, NT-Linear exhibits unstable classification performance due to over-compression of standard attention mechanisms. Compared to other baselines, NetMamba achieves improved memory efficiency primarily by customizing GPU operators that minimize the storage of extensive intermediate states and conduct recomputation during the backward pass.
@@
NetMamba (default) 0.9094 0.9096 0.9301 0.9305 0.9928 0.9929 0.9986 0.9986 0.9805 0.9806 0.9960 0.9957
w/ Bidirectional Mamba1 0.9012 0.9015 0.9213 0.9191 0.9974 0.9974 0.9966 0.9966 0.9704 0.9706 0.9951 0.9951
w/ Cascading Mamba1 0.8194 0.8233 0.9015 0.8998 0.9687 0.9687 0.9952 0.9952 0.9320 0.9318 0.9852 0.9845
w/ Vanilla Transformer2 0.8836 0.8837 0.9058 0.9045 0.9938 0.9939 0.9973 0.9973 0.9632 0.9633 0.9954 0.9952
w/ Linear Transformer2 0.6413 0.6857 0.4226 0.4717 0.8447 0.8558 0.8471 0.8708 0.7023 0.7361 0.7502 0.7742
w/o Position Embedding 0.9091 0.9092 0.9103 0.9081 0.9872 0.9872 0.9979 0.9979 0.9588 0.9586 0.9920 0.9920
w/o Pre-training 0.8868 0.8874 0.9151 0.9142 0.9882 0.9882 0.9966 0.9966 0.9335 0.9346 0.9904 0.9904
w/o Header 0.5814 0.6626 0.7750 0.8124 0.5597 0.5318 0.8340 0.8434 0.5058 0.4394 0.5085 0.5672
w/o Payload 0.9010 0.9004 0.9365 0.9362 0.9856 0.9856 1.0000 1.0000 0.9747 0.9747 0.9944 0.9945
w/ Patch Splitting3 0.8857 0.8863 0.9125 0.9103 0.9892 0.9892 0.9973 0.9973 0.9617 0.9619 0.9889 0.9890
@@ VII-D1 Model-level Ablation
Initially, we replaced all unidirectional Mamba blocks in NetMamba with bidirectional ones used by [12]. The experimental results revealed a slight performance decline across all datasets except for CICIoT2022. This suggests that unidirectional Mamba is well-suited for processing network traffic data, given that packets are transmitted sequentially and earlier packets possess limited information about subsequent ones. Moreover, incorporating bidirectional or even omnidirectional Mamba blocks introduces additional computational and memory overheads due to extra scan passes, ultimately reducing efficiency. Thus, unidirectional Mamba stands out as the preferable choice.
@@
To further investigate the expressiveness of the Mamba architecture, we evaluated two Transformer-based ablation variants, NT-Vanilla and NT-Linear. The results show that NT-Vanilla performs slightly worse than NetMamba across all datasets except for CICIoT2022. This indicates that the linear-time Mamba model is well suited for capturing sequential traffic data using our proposed representation scheme. Moreover, NT-Linear, due to information loss from the over-compression of its attention mechanisms, exhibits unstable performance and falls significantly behind NetMamba on three datasets.
While positional information is inherently preserved in sequence models such as Mamba, eliminating explicit positional embedding still results in a reduction in accuracy ranging from 0.03% to 2.17% across all datasets. This suggests that reinforced positional information aids the model in capturing correlations within sequential traffic data.
The pre-training process is designed to capture general traffic understanding from extensive unlabeled data. When compared to the non-pre-trained counterpart, pre-trained NetMamba demonstrates accuracy improvements ranging from 0.20% to 4.70%, affirming the effectiveness of our MAE-based pre-training task.